I would really like it if you guys can also help me with this application problem.
It reads;

An inlaid checkerboard has a mohogany border 2 inches wide. The area of the border is 176 square inches. Find the length of the board.
Thank you for your time.
C. Johnson

Answer by checkley75(3666)

2*4(x+2)=176
8(x+2)=176
8x+16=176
8x=176-16
8x=160
x=160/8
x=20 inches for the side of the checkerboard.
proof
20*20=400 sq. in.
24*24=576 sq. in.
576-400=176 sq. in. for the border.